OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = -x + 1, passing through the point (4,1)? A. x + y = 5 B. x - y = -5 C. -x - y = 5 D. x + y = -5 ***Please explain! :) THx!! :D

jimthompson5910 (jim_thompson5910):

what's the slope of y = -x + 1

OpenStudy (anonymous):

-1 right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

the slope of any parallel line will also have a slope of -1

jimthompson5910 (jim_thompson5910):

so m = -1 for this new line

jimthompson5910 (jim_thompson5910):

now plug this along with (4,1) into y = mx+b y = mx+b y = -1x+b ... Plug in m = -1 1 = -1(4)+b ... Plug in (4,1)...or replace x with 4 and y with 1 Now solve for b

OpenStudy (anonymous):

1=-4+b 5=b

OpenStudy (anonymous):

is that right??

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so the equation of the parallel line is y = -x+5

jimthompson5910 (jim_thompson5910):

now convert to standard form

OpenStudy (anonymous):

 is that right?